o at 22 oto stam at ihe spec

please show work
Answer #1

If we are converting water into the steam, then heat goes in the following steps

1 to raise the temperature of water from 22*C to 100*C

2 To convert the water at 100*C to steam at 100*C(Phase change)

3 To raise the temreature of steam from 100*C to 120*C

So total amount of heat energy required is the algebric sum of energy reuired in all three steps. so calculate the energy in every steps, so lets start with 1.

h=mc_{v}\Delta T

where temperature difference iss 100-22= 78 and specific heat Cv is 4.184 and mass is 250g

h_{1}=0.200g*4186*78=65301.6j= 65.301Kj

2. to change in phase we required some heat energy to 100*C water which is given by

h_{2}= m\Delta H_{Vap}

h_{2}= 0.200g*22.6*10^{5}=4.52*10^{5}j=452Kj

3. To convert the 100*C waterinto 120*C steam is given by(specific heat of steam is taken 1890, sometimes 1860) so we are taking 1890

h_{3}= mC_{steam}*\Delta T

h_{3}= 0.200*1890*20=7560j= 7.56Kj

hence total heat energy required is:

h_{1}+h_{2}+h_{3}= +7.56Kjh_{1}+h_{2}+h_{3}= 65.30Kj+452Kj+7.56Kj=524.56Kj

So total heat energy required is 524.56Kj